# Robust High-Fidelity Quantum Entanglement Distribution over Large-Scale Metropolitan Fiber Networks with Co-propagating Classical Signals

### v1

@@ III.1 CHSH over Local & Urban Distances
To characterize our quantum network operation, we measure the CHSH inequality violation and the fidelity of our quantum states relative to the |$\ket{\Phi_{+}}$ Bell state. We establish a baseline for the performance of the entanglement distribution by performing tests both locally and over a 60 km fiber link without APC or path routing. The local CHSH test yields an $S$-value of $S=2.71\pm 0.01$ and a quantum fidelity of $96.84\%\leq F_{Q}\leq 98.31\%$ (Figure 4 a). The same measurement was performed through the switch and over 60 km resulting in a $S$-value of $S=2.60\pm 0.04$ and $95.31\%\leq F_{Q}\leq 96.52\%$ (Figure 4 b). The somewhat worse performance of the 60 km loop is due to the increased SRC pair rate necessary to overcome the fiber losses, as explained in Section II.2.
Flexible path routing is necessary for a truly hybrid network operation that is in accordance with current classical networks run by telecom operators. To enable routing, we introduce automated path-switching in our setup. The APC re-aligns polarization after each path change, ensuring high entanglement fidelity regardless of the route. To confirm that the additional hardware does not negatively affect the results, we repeat the measurement locally with all the devices in place. An $S$-value of $S=2.72\pm 0.01$ is measured, consistent with the initial local results (Figure  4 c). After switching to the 60 km path and running a compensation routine, we measure $S=2.65\pm 0.10$ and quantum fidelity of $96.34\%\leq\text{F}_{Q}\leq 98.38\%$ (Figure  4 d), once again lower due to higher pair rates. These results demonstrate that automated path-switching and polarization compensation function as intended, eliminating the need for manual adjustments and enabling stable quantum correlations over independent fiber paths.
Reference Location Setting Fiber (km) Fidelity (%) Pair Rate (pairs/s) Quantum $\lambda$ (nm) Classical $\lambda$ (nm) Duration This work Berlin, DE Urban $10^{-2}$ to 82 99-85 107- $5\times 10^{2}$ 795, 1324 1560.61 days (ea.) [22] Chicago, US Laboratory 30.2 $>90$ 1290-1310 1547 - [23] Geneva, CH Laboratory 50 $\sim 80$ $\sim 10^{1}$ 1310 1534 hours [24] Singapore, SG Laboratory 50 92.5-97 $10^{4}$ 586, 1310 - - [29] Brooklyn, US Urban 34 99 $5\times 10^{5}$ 795, 1324 - 15 days [19] Stuttgart, DE Urban 35.8 94.5 $\sim 10^{3}$ 780, 1515 - days [20] Chicago, US Urban 47.9 $\sim 99$ $\sim 10^{2}$ 1280-1318 1549-1565 - [21] Shandong, CN Urban 66 N/A N/A 1310 1490, 1538 -
Table 1: Comparison of performance metrics of O-band quantum entanglement distribution or comparable testbeds [^2] To the best of the author’s knowledge these represent the state-of-the-art of O-band or bichromatic photon-pair quantum entanglement distribution. We distinguish two types of experiments: fiber spool-based in-lab setups (incl. between nearby buildings supplemented with spools) (laboratory); real-world deployments using installed fiber (urban). If a metric was unavailable, we used “-”; if not relevant, we used “N/A”.
@@ III.2 Sustained Automated Path Switching
A second benchmark to establish BearlinQ as a mature telecom network is the demonstration of long-term reliable operation. This is shown by continuously operating the system over multiple days with automated switching across available paths. To assess the robustness of the adaptive routing, periodic CHSH and fidelity measurements are performed while alternating the photon route between the local T-lab path and the long-distance fibers every 30 minutes. Each time the fiber path is switched, the APC’s compensation routine is triggered. A time diagram of the routine is shown in Figure 5 a. With a more advanced optical switch and access to additional fiber paths, this approach could scale to route polarization qubits from multiple entanglement sources to various nodes across a metropolitan-scale quantum network.
@@
We run long-term measurements for 15 km, 30 km, 60 km, and 82 km paths over durations ranging from 16 to 65 hours. The 15 km, 30 km, and 60 km measurements are performed in parallel with periodic switching between configurations and active compensation dynamically adapting to varying fiber conditions. Figure 5 b shows the classical APC-calculated start fidelity (pre-compensation) and end fidelity (post-compensation) for each switching event to illustrate the effectiveness of polarization compensation. We see that different fiber paths require distinct polarization transformations, resulting in variable compensation times. This comes at the cost of network downtime, which ranges from tens of milliseconds to a few seconds (average compensation time of 511 ms across all paths of Figure 5). The $95^{\text{th}}$ percentile compensation times for a change from local to 60 km path and vice versa are $9.5$ s and $250$ ms, respectively. Figure  5 c, d, e, and f and Figure 6 show the calculated $S$-values and quantum fidelity bounds for each fiber length. We observe sustained quantum correlations across all configurations, with an average $S=2.66\pm 0.04$ and quantum fidelity bounds between $90\%\leq F_{Q}\leq 99\%$ whenever measuring on the local path. For the 15 km, 30 km, 60 km loops, we measure average CHSH parameters of $S=2.53\pm 0.05$, $S=2.61\pm 0.06$, and $S=2.61\pm 0.09$, respectively. On the 15 km path, the quantum fidelity lower bound exceeded $90\%$ during 75% of the runtime, while on the 30 km and 60 km fibers this threshold was upheld for 94% and 79% of the runtime, respectively. All paths exceed $85\%$ quantum fidelity lower bound $100\%$ of the time, except the 60 km path which does 98.5% of the time. Given the necessity of longer integration times, the 82 km data was taken without path switching. In this longest fiber path, we measure $S=2.58\pm 0.03$ and a fidelity $89.5\%\leq F_{Q}\leq 98.4\%$. All these measurements highlight robust long-term quantum operations over all fiber routes.
@@
Figure 6: Entanglement distribution over 82 km. The results confirm the system’s ability to preserve strong quantum correlations despite high losses ($-42$ dB) and uncontrolled environmental perturbations introduced by the fibers. The integration time for each waveplate configuration in the CHSH and fidelity measurements was 60 s.
@@ III.3 Operational Metrics
Network operators rely on a set of well-defined key performance indicators (KPIs) to assess critical daily services. In conventional telecom networks, these include high uptime (¿99.99%), non-interference between channels, and predictable signal integrity. Despite it being particularly challenging, translating these standards to the quantum domain is necessary to establish commercial hybrid quantum-classical networks. Achieving telecom-grade reliability in the quantum domain requires active stabilization, real-time diagnostics, and extremely low-noise operation, none of which are trivial with entangled photons. By addressing these issues, our system takes a step toward bridging the gap between quantum networks as a research tool and as real-world infrastructure.
BearlinQ meets telecom-inspired KPIs by demonstrating uninterrupted entanglement distribution while routing across metropolitan infrastructure under real-life conditions. Operator-aligned metrics such as low downtime, minimal impact on classical traffic, high rates, and high-fidelity qubit transmission are achieved. In this context, downtime is defined as periods when $F_{Q}$ falls below $85\%$ or the system requires calibration or maintenance. Automatic stabilization ensures near carrier-grade reliability, as demonstrated by $1.5\%$ downtime on the 60 km path while maintaining quantum fidelity $F_{Q}>85\%$ (Figure 5 d). Quantum fidelity was primarily limited by the high entanglement generation rate of the SRC and residual detuning in the APC (see Supp. Mat. IV.2). An additional minor reduction of $-0.3\%$ was attributed to polarization-dependent loss (PDL) totaling $\sim 0.9$ dB, excluding long-haul links, mainly arising from the optical switch and WDM components.
Beyond classically-inspired KPIs, hybrid networks have quantum-specific constraint. As explained in Sec.  II.2, noise in the quantum channel is one of them. To minimize unwanted photons at the quantum wavelengths, our network employs a quantum signal on the 1310 nm CWDM channel mixed using OADMs with classical data on C-band DWDM channels at 1551.32 nm (Ch 27, +6 dBm) and 1560.61 nm (Ch 13, +4 dBm). The results in Figure 7 demonstrate low noise levels in the 60 km path configuration, enabling effective copropagation.
Figure 7: Noise pollution analysis on the 60 km link. Average of correlation peaks for all waveplate configurations in a full CHSH measurement with (magenta) and without (blue) co-propagating classical light. The uncorrelated background noise, primarily from detector dark counts and the SRC’s thermal emission, yields SNRs of 95 (classical off) and 90 (classical on). The classical channel introduces $\sim 30$ additional counts/s on average.
@@
Practical quantum networking also demands useful pair-rates. Instrumental for this result was the high SRC pair-generation rate of $\geq 10^{7}$ pairs/s. Our deployed system consistently delivers coincidence rates ranging from $5\times 10^{2}-10^{7}$ entangled pairs/s after losses, depending on the selected fiber path. Even at attenuation levels of up to $-42$dB (82 km path), our network had $\geq 10^{2}$ pairs/s, showcasing a resilient platform for real-world useful quantum networking. However, in the absence of ultralow-loss fibers or quantum repeaters, the attenuation of the O-band fiber at $\sim$0.32 dB/km limits practical point-to-point deployments to distances of $\lesssim$100 km. Prior O-band demonstrations relied on fiber spools in laboratory settings [23, 24] or short campus loops [25]. Here, we exceed those benchmarks using 82 km of urban fiber despite added insertion losses from the adaptive routing and non-ideal deployed fibers. Table 1 summarizes key performance metrics and highlights how this deployment establishes a new operational benchmark for metropolitan quantum networks.
@@ IV Conclusion
Practical quantum networking requires critical components including bright, deployable entanglement sources producing high-fidelity photon pairs, rapid reference frame stabilization, and low-loss routing. Here, we addressed these challenges and successfully deployed BearlinQ, a quantum testbed in the metropolitan fiber network of Deutsche Telekom in Berlin. Using commercially available devices and telecom-standard optical add-drop multiplexing hardware, we demonstrated uninterrupted high-rate entanglement distribution through dynamically selectable fiber paths ranging from 10 m to 82 km with quantum signals in the telecom 1324 nm O-band seamlessly coexisting with classical bidirectional C-band traffic. Using active polarization stabilization, our system maintained high entanglement fidelity, with lower and upper bounds ranging from 85% to 99% and with S values between 2.36 and 2.74, indicating strong quantum non-locality, with near telecom-grade reliability ($\sim 1$% downtime) over extended durations, all without requiring dedicated fiber or modifications to existing telecom infrastructure. Our deployment also validates the feasibility of quantum routing within existing telecom infrastructure, demonstrating telecom operational reliability requirements over typical metropolitan distances.
By leveraging co-existence of usable quantum photons with standard classical telecom traffic, this work sets operational benchmarks necessary for commercially viable quantum services, eliminating costs and complexity of new fiber installations. These results highlight that quantum networks can be integrated with existing assets and are ready to transition from lab-based demonstrations to reliable deployments. By embracing these innovations, operators can establish themselves as the backbone of quantum-enabled networks, paving the way toward commercially viable quantum services for quantum communications, distributed quantum computing, and quantum sensing applications integrated into existing urban infrastructure.
@@ IV.1 Performance Optimizations
CHSH consistency: The calibration procedure begins by inserting perpendicularly aligned polarizers in the SRC to preselect the polarization state |$\ket{H}$ or |$\ket{V}$. Motorized $\lambda/2$-waveplates are then aligned with their slow/fast axes set to the zero angle position, which is then stored as a home reference. Next, manual polarization controllers before the measurement station are adjusted to minimize single-photon counts, thus defining the reflected basis of the PBS’s as the reference basis of measurement. Subsequently, the waveplates are rotated to the diagonal basis where LCR voltages are then fine-tuned to minimize coincidences. These steps could be automated using programmable fiber paddles, or the APC itself, eliminating user-induced variability and enabling more precise and repeatable calibrations than those presented herein.
Fidelity improvements: High-fidelity entanglement distribution relies on spectrally narrowband photons to minimize internal wavelength spread and associated polarization dispersion. However, quantum fidelity is fundamentally constrained when the APC is calibrated at a wavelength different from that of the quantum signal because fiber-induced polarization rotations are both wavelength-dependent and time-varying. Measurements in deployed telecom fibers show that the fiber’s polarization transformation varies significantly with wavelength. This wavelength dependence becomes pronounced over tens of kilometers, rendering a system calibrated at one wavelength inherently misaligned for others. When polarization compensation is performed at 1324 nm, high fidelity is maintained only for photons within a narrow spectral range (¡0.5 nm). As shown in Figure 8, even slight detuning (or a broad photon bandwidth) introduces measurable infidelities. Our data show that spectral mismatch consistently reduces Bell-state fidelity across all fiber configurations. This reduction was the high fiber attenuation limiting APC operation at the target wavelength due to insufficient optical power, leading to a quantum fidelity losses of up to 4%. This constraint can be easily mitigated in future implementations.
@@
Operational downtime: To minimize network interruptions, we plan to implement continuous monitoring of polarization drift by analyzing a tap on existing co-propagating C-band light with a 15XX nm APC-CMP. This unit would then take advantage of its embedded polarimeter to detect SOP discontinuities, which would be used to trigger the 13XX nm APC. This stabilization protocol eliminates the need to rely solely on routine classical fidelity checks in the time-multiplexed compensation scheme, thereby reducing network downtime.

### v2

@@ III.1 CHSH over Local & Urban Distances
To characterize our network, we measure the CHSH inequality violation and the quantum state fidelity relative to the |$\ket{\Phi_{+}}$ Bell state. We establish a baseline for the performance of the entanglement distribution both locally and over a 60 km fiber link without APC, nor path routing. The local CHSH test yields an $S$-value of $S=2.71\pm 0.01$ and a quantum fidelity of $96.84\%\leq F_{Q}\leq 98.31\%$ (Figure 4 a). The same measurement was performed adding the switch and over 60 km resulting in a $S$-value of $S=2.60\pm 0.04$ and $95.31\%\leq F_{Q}\leq 96.52\%$ (Figure 4 b). The fidelity reduction at 60 km stems from the increased pair generation rate required to overcome loss, which in turn degrades signal-idler correlations.
Flexible path routing is necessary for network operation in accordance with current classical networks and future use-cases. To enable routing, we introduce automated path-switching in our setup. The APC re-aligns polarization after each path change, ensuring high entanglement fidelity regardless of the route. To confirm that the additional hardware does not negatively affect the results, we repeat the measurement locally with all the devices in place. An $S$-value of $S=2.72\pm 0.01$ is measured, consistent with the initial local results (Figure  4 c). After switching to the 60 km path and running a compensation routine, we measure $S=2.65\pm 0.10$ and quantum fidelity of $96.34\%\leq\text{F}_{Q}\leq 98.38\%$ (Figure  4 d), once again lower due to higher pair rates. These results demonstrate that automated path-switching and polarization compensation function as intended, eliminating the need for manual adjustments and enabling stable quantum correlations over independent fiber paths.
The consistency of CHSH curves depends on the calibration procedure which begins by inserting perpendicularly aligned polarizers in the SRC to preselect the polarization state |$\ket{H}$ or |$\ket{V}$. Motorized $\lambda/2$-waveplates are then aligned with their slow/fast axes set to the zero-angle position, which is then stored as a zero-position reference. Next, manual polarization controllers before the measurement station are adjusted to minimize single-photon counts, thus defining the reflected basis of the PBS’s as the reference basis of measurement. Subsequently, the waveplates are rotated to the diagonal basis, where LCR voltages are then fine-tuned to minimize coincidences. These steps could be automated, eliminating user-induced variability and enabling more precise and repeatable calibrations than those presented herein.
@@ III.2 Sustained Automated Path Switching
A second benchmark to establish BearlinQ as a exemplary network is the demonstration of long-term reliable operation. This is shown by continuously operating the system for multiple days, including automated switching across the available paths when integration times allowed. To assess the robustness of the adaptive routing protocols, periodic CHSH and fidelity measurements are performed while alternating the photon route between the local T-lab path and the long-distance fibers. Each time the fiber path is switched, the APC compensation is triggered. A timing diagram of the routine is shown in Figure 5 a. With the addition of further switch ports, paths, and entanglement sources, this architecture could scale towards a multi-user quantum routing framework.
The core of our reliability tests are long-term measurements for 15 km, 30 km, and 60 km paths over durations ranging from 16 to 65 hours. All measurements are performed in parallel with periodic switching between configurations and active compensation dynamically adapting to varying fiber conditions (Figure 5).
@@
To illustrate the effectiveness of polarization compensation, Figure 5 b shows the classical APC-calculated start fidelity (pre-compensation) and end fidelity (post-compensation) for each switching event. We see that different fiber paths require distinct polarization transformations, resulting in variable compensation times and at the cost of network downtime, which ranges from tens of milliseconds to a few seconds (average compensation time of 511 ms across all paths of Figure 5). The $95^{\text{th}}$ percentile compensation times for a change from local to 60 km path and vice versa are 9.59.59.5 s and 250250250 ms, respectively.
Figure  5 c, d, e, and f and show the calculated $S$-values and quantum fidelity bounds for each fiber length employed in the path-switching experiment. We observe sustained quantum correlations across all configurations, with an average $S=2.66\pm 0.04$ and $90\%\leq F_{Q}\leq 99\%$ whenever measuring the local path. For the 15 km, 30 km, 60 km loops, we measure average CHSH parameters of $S=2.53\pm 0.05$, $S=2.61\pm 0.06$, and $S=2.61\pm 0.09$, respectively. On the 15 km path, the quantum fidelity lower bound exceeded 90% during 75% of the runtime, while on the 30 km and 60 km fibers this threshold was maintained for 95% and 79% of the runtime, respectively. All paths exceed the 85% quantum fidelity lower bound 100% of the time, except the 60 km path which does 98.5% of the time. Collectively, these results demonstrate robust long-term operations and the flexibility of our network.
Refer to captionRefer to caption
Figure 5: Sustained automated multi-path switching in a metropolitan network. a) Sequence of logic operations showing key timing intervals for CHSH measurement (blue, $t_{\mathrm{CHSH}}$), active polarization control (green, $t_{\mathrm{APC\,comp}}$), path switching (yellow, $T_{\mathrm{path\,switch}}$), and local acquisition (yellow, $t_{\mathrm{local}}$). Vertical dotted lines mark switching events. b) Start and end classical fidelity for all APC routines during the switch from BearlinQ to local fiber. Histograms at right show the distribution of final classical fidelities measured by the APC. c) Local long-term CHSH and quantum fidelity traces, each corresponding to a switch from one of the deployed fibers (see following panels); similar color tones indicate the same measurement run. d), e), f) Long-term CHSH and quantum fidelity over 15 km, 30 km, and 60 km, respectively. Error bars in c–f show $1\sigma$ uncertainties from Poissonian count statistics propagated through the CHSH $S$ calculation. In c–f, the red dashed line marks $S=2$. The integration time for each waveplate configuration was 1 s for the local, 15 km, and 30 km paths, and 2 s for the 60 km path.
@@
Figure 6: Entanglement distribution over 100 km through Berlin. These results demonstrate the system’s ability to maintain strong quantum correlations despite significant transmission losses and persistent uncontrolled environmental perturbations in the deployed fiber. This dataset was acquired without the path switching equipment included. Each waveplate setting in the CHSH and fidelity measurements was integrated for 30 s.
To test the limits of our testbed, we also run fidelity and CHSH measurements over 100 km. However, due to the need for longer integration times, this data was acquired without path switching nor DWDMs. Figure 6 shows the $S$-values and quantum fidelity bounds for the 100 km loop run continuously for 78 hours. In this experiment, the system achieved $S=2.58\pm 0.03$ and a fidelity between $80.5\%\leq F_{Q}\leq 98.4\%$.
@@ III.3 Operational Metrics
Network operators rely on a set of key performance indicators (KPIs) to evaluate service performance. In conventional telecom networks, these include uptimes exceeding $99\%9$, negligible interchannel interference, and consistent signal integrity. Although particularly challenging, translating these expectations into the quantum domain is essential for building practical hybrid quantum-classical networks. Achieving telecom-grade reliability in quantum systems requires active stabilization, real-time diagnostics, and ultra-low-noise operation, all nontrivial engineering feats. Our system addresses these challenges, taking a step toward transitioning quantum networks from experimental setups to real-world infrastructure. It demonstrates quasi-uninterrupted entanglement distribution over metropolitan-scale infrastructure under real operating conditions.
[…]
@@
Practical quantum networking also demands useful pair-rates. Instrumental for this result was the high SRC pair generation rate of $\geq 10^{7}$ pairs/s. Our deployed system consistently delivers coincidence rates ranging from $5\times 10^{2}-10^{7}$ entangled pairs/s after losses, depending on the selected fiber path. However, in the absence of ultralow-loss fibers or quantum repeaters, the attenuation of the O-band fiber at $\sim$0.31 dB/km limits practical point-to-point deployments to distances of $\sim$100 km. Previous O-band demonstrations relied on fiber spools in laboratory settings [24, 25] or short campus loops [26]. Here, we exceed those benchmarks by reaching a rate of $\geq 10^{2}$ pairs/s using up to 100 km of urban fiber (45-45- 45 dB attenuation), demonstrating a resilient platform for useful quantum networking in the real world.
Table 1 summarizes key performance metrics and highlights how this deployment establishes a new operational benchmark for metropolitan quantum networks.
[…]
Table 1: Comparison of performance metrics of O-band quantum entanglement distribution or comparable testbeds [^2] To our knowledge these represent the state-of-the-art of O-band or bichromatic photon-pair entanglement distribution. We distinguish two types of experiments: fiber spool-based in-lab setups or nearby buildings supplemented with spools (laboratory), and real-world deployments using installed fiber (urban). If a metric was unavailable, we used “-”; if not relevant, we used “N/A”.
Lastly, beyond classical-inspired KPIs, hybrid networks must also meet quantum-specific requirements, the most critical being low noise in the quantum measurements (see Sec.II.2). In our setup, quantum signals are transmitted in the 1310 nm CWDM band, while classical traffic co-propagates in the C-band using a DWDM channel at 1560.61 nm (Channel 21 at +4 dBm launch power), combined via OADMs. Unlike the entangled photons, the background noise and stray classical light don’t change significantly in intensity as the waveplate rotates before the polarizer, showing they’re mostly unpolarized. As shown in Figure 7, the 60 km path maintains similar noise levels when the quantum signal is distributed in the presence or absence of classical light, confirming the viability of co-propagation.
Figure 7: Noise pollution analysis on the 60 km link. Averaged correlation peaks across all waveplate settings in a full CHSH measurement, with (orange) and without (green) co-propagating classical light. The background noise, dominated by detector dark counts and thermal emission from the SRC, results in SNRs of 35 with classical ON, and 49 with classical OFF.
@@ IV Conclusion
Practical quantum networking requires critical components including bright, deployable entanglement sources producing high-fidelity photon pairs, rapid reference frame stabilization, and low-loss routing. Here, we addressed these challenges and successfully deployed BearlinQ, a quantum testbed in the metropolitan fiber network of Deutsche Telekom in Berlin.
Using commercially available devices and telecom-standard optical add-drop multiplexing hardware, we demonstrated uninterrupted high-rate entanglement distribution with quantum signals in the telecom 1324 nm O-band seamlessly coexisting with classical bidirectional C-band traffic along dynamically selectable fiber distances ranging from 10 m to 60 km. With active polarization stabilization, our system maintained high entanglement fidelity, with lower and upper bounds ranging from 85% to 99% and with $S$ values between 2.36 and 2.74, indicating strong quantum nonlocality, with near telecom-grade reliability (<1.5<1.5< 1.5% downtime) over extended durations, all without requiring dedicated fiber or modifications to existing telecom infrastructure. Our deployment also validates the feasibility of quantum routing within existing telecom infrastructure, demonstrating telecom operational reliability requirements over typical metropolitan distances.
By leveraging the co-existence of usable quantum photons with standard classical telecom traffic, this work sets operational benchmarks necessary for commercially viable quantum entanglement-based services, eliminating the costs and complexity of new fiber installations. These results highlight that quantum networks can be integrated with existing assets and are ready to transition from lab-based demonstrations to reliable deployments. By embracing these innovations, operators can establish themselves as the backbone of quantum-enabled networks, laying the technical foundation for commercially viable quantum services for quantum communications, distributed quantum computing, and quantum sensing applications integrated into existing urban infrastructure.
@@ IV.1 Performance Optimizations
Fidelity improvements: High-fidelity entanglement distribution relies on spectrally narrowband photons to minimize internal wavelength spread and associated polarization dispersion. However, quantum fidelity is fundamentally constrained when the APC is calibrated at a wavelength different from that of the quantum signal because fiber-induced polarization rotations are both wavelength-dependent and time-varying. Measurements in deployed telecom fibers show that the fiber’s polarization transformation varies significantly with wavelength. This wavelength dependence becomes pronounced over tens of kilometers, rendering a system calibrated at one wavelength inherently misaligned for others. When polarization compensation is performed at 1324 nm, high fidelity is maintained only for photons within a narrow spectral range (¡0.5 nm). As shown in Figure 8, even slight detuning (or a broad photon bandwidth) introduces measurable infidelities. Our data show that spectral mismatch consistently reduces Bell-state fidelity across all fiber configurations. This reduction was due to high fiber attenuation that limited the operation of APC at the target wavelength due to insufficient optical power, leading to quantum fidelity losses of up to 5%. This constraint can be mitigated.